OpenStudy (blondey_315):

Which ratio is NOT equal to 16/24?

OpenStudy (anonymous):

where is the choice?

OpenStudy (blondey_315):

oh sorry! 4/8 2/3 4/6 12/18

OpenStudy (anonymous):

16/24 = 4/6 = 2/3

OpenStudy (anonymous):

4/8

OpenStudy (anonymous):

so the answer is 4/8

OpenStudy (anonymous):

Would you like an explanation?

OpenStudy (anonymous):

^

OpenStudy (anonymous):

Wait, I wonder, what grade work is this again?

OpenStudy (anonymous):

All the other answer's are the same ratio, just divide them by 4 and you will always get 2/3, but with 4/8 you get 1/2!

OpenStudy (anonymous):

@kmcllb : you need to explain NOT just only give the answer dude

OpenStudy (anonymous):

I did. :)

OpenStudy (anonymous):

good

OpenStudy (anonymous):

@gerryliyana Will you help me with my equation?

OpenStudy (anonymous):

of course, go ahead

OpenStudy (anonymous):

Which is an equation of the line with slope -3/4 passing through the point (-,2,-10)? A. y = -3/4x -23 B. y = -3/4x = 19/2 C. y = -3/4x - 10 D. 3x - 4y = 23 Please give me step by step, thank you!

OpenStudy (anonymous):

@gerryliyana

OpenStudy (anonymous):

hold on

OpenStudy (anonymous):

@kmcllb If we have a point ( x1 , y1 ), and a slope, m, here's the formula we use to find the equation of a line: y - y1 = m(x-x1) It's called the point-slope formula! Let's find the equation of the line that passes through the point (-,2,-10) with a slope of -3/4 x1 = -2 y1 = -10 m = -3/4 So y - y1 = m(x - x1) y - (-10) = (-3/4) * ( x - (-2)) y + 10 = (-3/4) * (x+2) y = (-3/4) * (x+2) - 10 y = -3/4x - 10 y - (-10) = m (x -)

OpenStudy (anonymous):

the answer is y = -3/4x - 10